A bucket of water with a total weight of 50 Newtons is lifted at constant velocity up a 10 meter deep well.If it takes 20 seconds to aise the bucket this distance, the power required to lift the bucket is:

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Power = force * distance/ time = 50*10/20= 25 watts

Regards
Vikas TU
4 years ago
Dear student 
Power = F.V 
Velocity = distance/time 
Velocity = 10/20 = 0.5 m/s 
P = 50 * 0.5 = 25 Watt
Good Luck